b1az3:

Solve for x. the square root of 3x-1=7 the square root of 3x+1=7 What is the root? If there is no root, choose none. 16 none 12 16 2/3

Shadow:

\[\sqrt (3x - 1)= 7\]

Shadow:

Is this what it looks like?

Shadow:

I'm just asking if the sqrt is just over the 3x or also extends to 1

b1az3:

Shadow:

Okay that works.

Shadow:

\[\sqrt (3x - 1)= 7\] Square both sides \[3x - 1 = 49\] Add 1 to both sides \[3x = 50\] Divide both sides by 3 \[x = 16.67\]

b1az3:

16 none 12 16 2/3 these are the only options

Shadow:

\[\sqrt (3x + 1) = 7\] \[3x + 1 = 49\] \[3x = 48\] \[x = 16\] Using the same logic

Shadow:

Yeah .67 is 2/3